OpenStudy (anonymous):

solve: 7/13*x=56/117

OpenStudy (anonymous):

just transpose

OpenStudy (anonymous):

??

OpenStudy (anonymous):

7/13*x=56/117 x= 56/117 * 13/7 =8/9

OpenStudy (anonymous):

ok thanks